OpenStudy (anonymous):

can someone help me please with this problem :( no one answer :( A ladder is leaning up against an apartment building to help people get out of the building during a fire. The base of the ladder is 7 feet from the building. The ladder is 1 foot longer than the height it reaches on the building. What is the length of the ladder? A. 12 feet B. 13 feet C. 24 feet D. 25 feet

OpenStudy (anonymous):

c

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @Bangerz_14 Post your question in Math to get better help :) http://www.openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

Thank you soo much @heather2292 I'll give u a medal :)

OpenStudy (anonymous):

 So, do you know how to solve that?

OpenStudy (brunsonni):

You use Pythagorean theorem for this. A^2 + B^2 = C^2 This is for right triangles, C is the hypotenuse, the diagonal line. If it's _ feet from the house, you fill in _^2 +B^2 = C^2 (I'm on a mobile device so I can't go back to see the problem lol sorry) whatever the other side of the right angle is, fill it in for B. Square the two numbers and add the products. So if it's 3^2 + 4^2 = C^2 you get 9+16=C^2 which is 25=C^2 so finally you find the square root of 25 which is 5, which is your hypotenuse.

OpenStudy (brunsonni):

So? Someone else could be looking for help with this and that might help them.

OpenStudy (anonymous):

belive its c